OpenStudy (anonymous):

cos65°cos5°+sin65°sin5°

OpenStudy (anonymous):

How would I solve this using the sum/diff formula double angle formula or 1/2 angle formula

OpenStudy (nurali):

cos(a-b)=cosacosb+sinasinb cos(65°)cos(5°) + sin(65°)sin(5°) =cos((65-5)°) =cos(60°) =1/2

OpenStudy (solomonzelman):

1=cos(0)=cos(65-65)=cos65°cos5°+sin65°sin5°

OpenStudy (solomonzelman):

If I remember correctly that cos(0) equals 1.

OpenStudy (anonymous):

hold up im absorbing this

OpenStudy (solomonzelman):

I messed up, sorry, I though it is all 65...

OpenStudy (solomonzelman):

@Nurali is right :)

OpenStudy (anonymous):

awesome. thanks for helping

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

thanks nurali

OpenStudy (anonymous):

so we just reversed the formula haha